Problem 1-19 Traditional and Contribution Format Income Statements [LO1-6] Todrick Company is a merchandiser that reported the following information based on 1,000 units sold Sales Beginning merchandise inventory Purchases Ending merchandise inventory Fixed selling expense Fixed administrative expense Variable selling expense Variable administrative expense Contribution margin Net operating income $ 210,000 $ 14,000 $ 140,000 $ 7,000 $ 8,400 $ 10,500 $ 42,000 $ 12,600 Required: 1. Prepare a contribution format income statement. 2. Prepare a traditional format income statement. 3. Calculate the selling price per unit. 4. Calculate the variable cost per unit. 5. Calculate the contribution margin per unit. 6. Which income statement format (traditional format or contribution format) would be more useful to managers in estimating how net operating income will change in responses to changes in unit sales?Complete this question by entering your answers in the tabs below. Req 1 Req 2 Req 3 to 5 Req 6 Prepare a contribution format income statement Todrick Company Contribution Format Income Statement Variable expenses Fixed expenses:Complete this question by entering your answers in the tabs below Req 2 Reg 3 to Reg 6 Req 1 Prepare a traditional format income statement. Todrick Company Traditional Income Statement Selling and administrative expenses:Req 1 Req 2 Req 3 to5 Req 6 3. Calculate the selling price per unit. 4. Calculate the variable cost per unit. 5. Calculate the contribution margin per unit. 3. |Selling price per unit 4. Variable cost per unit 5. Contribution margin per unit
Answer #1

SOLUTION

1. Contribution format income statement-

Particulars Amount ($) Amount ($)
Sales 210,000
Variable expenses:
Cost of goods sold (14,000+140,000-7,000) 147,000
Selling expenses 10,500
  Administrative expense (168,000-147,000-10,500) 10,500
Total variable expenses (210,000-42,000) 168,000
Contribution margin 42,000
Fixed expenses:
  Selling expenses (29,400-8,400) 21,000
  Administrative expense 8,400
Total fixed expenses (42,000-12,600) 29,400
Net operating income 12,600

2. Traditional format income statement-

Particulars Amount ($) Amount ($)
Sales 210,000
Less: Cost of goods sold 147,000
Gross profit 63,000
Expenses:
Variable selling expense 10,500
Variable administrative expenses 10,500
Fixed selling expense 8,400
Fixed administrative expenses 21,000 50,400
Net operating income 12,600

3. Selling price per unit = Sales / Units

= $210,000 / 1,000 units = $210

4. Variable cost per unit = Variable expenses / Units

= $168,000 / 1,000 units = $168

5. Contribution margin per unit = Contribution margin / Units

= $42,000 / 1,000 units = $42

6. Contribution format income statement would be better to estimate change in operating income due to change in sales unit.